# Med-R2: Crafting Trustworthy LLM Physicians through Retrieval and Reasoning of Evidence-Based Medicine

### v1

@@ Abstract
In recent years, Large Language Models (LLMs) have exhibited remarkable capabilities in clinical scenarios. However, despite their potential, existing works face challenges when applying LLMs to medical settings. Strategies relying on training with medical datasets are highly cost-intensive and may suffer from outdated training data. Leveraging external knowledge bases is a suitable alternative, yet it faces obstacles such as limited retrieval precision and poor effectiveness in answer extraction. These issues collectively prevent LLMs from demonstrating the expected level of proficiency in mastering medical expertise. To address these challenges, we introduce Med-R2, a novel LLM physician framework that adheres to the Evidence-Based Medicine (EBM) process, efficiently integrating retrieval mechanisms as well as the selection and reasoning processes of evidence, thereby enhancing the problem-solving capabilities of LLMs in healthcare scenarios and fostering a trustworthy LLM physician. Our comprehensive experiments indicate that Med-R2 achieves a 14.87% improvement over vanilla RAG methods and even a 3.59% enhancement compared to fine-tuning strategies, without incurring additional training costs. Code and datasets are available at https://github.com/8023looker/Med-RR.
Med-R2: Crafting Trustworthy LLM Physicians through Retrieval and Reasoning of Evidence-Based Medicine
Keer Lu, Zheng Liang§, Da Pan§, Shusen Zhang§, Xin Wu§, Weipeng Chen§, Zenan Zhou§ Guosheng Dong§, Bin Cui, Wentao Zhang Center for Data Science, Academy for Advanced Interdisciplinary Studies, Peking University, School of CS & Key Lab of High Confidence Software Technologies (MOE), Peking University, §Baichuan Inc. keer.lu@stu.pku.edu.cn, dongguosheng@baichuan-inc.com, {bin.cui, wentao.zhang}@pku.edu.cn
@@ 1 Introduction
In recent years, Large Language Models (LLMs) have emerged as pivotal tools in the medical domain, redefining the contours of healthcare practice and research Abd-Alrazaq et al. (2023); Yang et al. (2023); Thirunavukarasu et al. (2023). Their ability to process and understand vast amounts of unstructured medical data positions them at the forefront of medical research Clusmann et al. (2023); Mumtaz et al. (2024), clinical decision-making Hager et al. (2024); Kim et al. (2024), and patient care Busch et al. (2024); Tripathi et al. (2024).
@@
However, despite their remarkable capabilities, LLMs encounter several challenges when applied specifically to healthcare settings:
C1: Inefficiency in Knowledge Acquisition. Existing approaches predominantly rely on pretraining and fine-tuning LLMs with datasets from the medical domain, the process of which is inherently compute-intensive and resource-demanding, especially as model sizes increase Singhal et al. (2023b); Luo et al. (2022). Moreover, training with outdated datasets can result in a lack of highly specialized expertise, which can lead to suboptimal clinical recommendations or misinform healthcare professionals He et al. (2022); Kandpal et al. (2023).
C2: Limited Precision for Medical Retrieval. Compared to domain-specific training, Retrieval-Augmented Generation (RAG) systems provide a cost-efficient solution, leveraging the external knowledge base to enhance content generation Lewis et al. (2020). The retrieval quality is critical in RAG systems, where inaccuracy or misinformation can heavily influence the effectiveness of LLMs’ augmentation. While some efforts Wang et al. (2024); Jeong et al. (2024); Long et al. (2024) have been made to improve the retrieval precision, they neglect the specific and highly professional nature of medical knowledge, where tailored retrieval enhancement for distinct medical scenarios remains insufficiently explored.
C3: Low Effectiveness in Answer Extraction. Considering the constraints imposed by the models’ context window length, it is essential to critically appraise the retrieved medical evidence for its validity, impact, as well as applicability, and integrate the most pertinent ones with existing clinical expertise for problem-solving Case and Swanson (1998); Guyatt et al. (1992). Nonetheless, current studies fail to develop targeted answer extraction methods tailored for healthcare scenarios, where the nuanced evaluation of evidence hierarchies and intricate reasoning are required Sackett et al. (1996).
@@
  • Challenges in Medical Scenarios. Through conducting a quantitative analysis of strategies aimed at enhancing models’ medical capabilities, we underscore the challenges prevalent in healthcare scenarios, including high computational consumption as well as poor efficiency in knowledge retrieval and extraction (Figure 1).
  • LLM Physician Framework. We present Med-R2, a novel LLM physician that integrates the Evidence-Based Medicine (EBM) principles for clinical problem-solving, bolstering models’ performances within medical contexts (Section 3).
  • Performance and Effectiveness. Our comprehensive experiments indicate that Med-R2  achieves a 14.87% improvement over the vanilla RAG methods, and even a 3.59% enhancement compared to the fine-tuning strategies without additional training expenses (Section 4 and Section 5).
2 Background and Related Work
Evidence-Based Medicine (EBM)  EBM refers to the application of the best available research to healthcare, which requires evidence integration with clinical expertise and patient values Sackett et al. (1996); Guyatt et al. (1992); Sackett (1997). Clinical questions can be categorized into several types, including diagnosis, therapy, prognosis, etiology, prevention, cost, etc. Case and Swanson (1998). Each category intersects with EBM principles by emphasizing the systematic collection, evaluation, and application of the best retrieved evidence to inform medical decision-making Guyatt et al. (2000), detailed in Appendix A.
@@ 2 Background and Related Work
LLMs for Medical Domain  As the application of LLMs expands, their deployment in the medical domain has become a widely discussed topic Clusmann et al. (2023); Zeng et al. (2020). Recent studies have concentrated on the direct use of real or synthetic medical data for the pretraining or fine-tuning of LLMs Thirunavukarasu et al. (2023); Singhal et al. (2023a). Prominent open-source milestones include ChatDoctor Li et al. (2023) which integrates real-world doctor-patient communication data for training, PMC-LLaMA Wu et al. (2024) pretrained on 4.9 million medical literature records, and MEDITRON Chen et al. (2023), a scaling series of medical pretrained models. However, such extensive training can be computationally intensive. In contrast, Retrieval-Augmented Generation (RAG) systems offer a more efficient alternative, achieving comparable results with reduced training costs and enhancing the model’s precision in locating and leveraging knowledge.
Retrieval-Augmented Generation (RAG)  The concept of RAG Lewis et al. (2020) was introduced as a powerful framework for integrating external knowledge into natural language generation tasks, enhancing the accuracy and relevance of generated outputs across various domains Gao et al. (2023); Zhao et al. (2024). In the medical field, RAG has been widely used to improve LLMs’ analytical performances by utilizing external medical knowledge from sources such as medical papers, textbooks, guidelines, and entries Jin et al. (2023); Zakka et al. (2024); Xiong et al. (2024). However, while there has been efforts dedicated to optimizing the individual components of RAG pipelines Wang et al. (2024); Jeong et al. (2024), research that integrates the unique characteristics and requirements of the medical domain remains in its infancy. In this study, we incorporate the principles of Evidence-Based Medicine (EBM) into medical RAG systems to better address the special demands of healthcare.
@@ 3.1 Question Formulation
The clinical question formulation stage consists of two components: query classifier and query reformulator. The query classification encompasses two dimensions: Evidence-Based Medicine (EBM) categories and general natural language question types. Specifically, we have delineated six distinct EBM categories and twelve general question categories, as illustrated in Figure 2. The application of this classification scheme to the MedQA-USMLE dataset yielded the categorized results presented in Figure 3. For the classification task, we employed Qwen2.5-72B-Instruct [^1] https://huggingface.co/Qwen/Qwen2.5-72B-Instruct as our classifier. We perform domain-specific reformulations of the original queries based on their respective classes of the EBM categories to align with the professional context. Meanwhile, the general question categories are utilized as one of the criteria to rerank retrieved documents, thereby prioritizing those that best match the current query’s intent and document type preferences described in Figure 8. Further details regarding the method and implementation can be found in Appendix B.
@@ 3.2.1 Evidence Retriever
Knowledge Corpus  To construct a more comprehensive medical knowledge base adaptable to diverse healthcare scenarios, we have amassed a collection of both open-source and in-house medical data to build our retrieval corpus. The final medical knowledge base employed for retrieval comprises four distinct types of resources: academic papers, entries, books, and guidelines, with details of the statistics depicted in Table 1. For resources with extensive content such as academic papers, books, and guidelines, we first perform content segmentation with a threshold set at 10,000 tokens. We prioritize dividing the content based on natural chapters. If natural chapters cannot be identified or exceed the threshold, we resort to truncation according to the predefined limit.
Source Type # Volume
Academic Papers 6,000,000
Entries 470,000
Books 10,000
Guidelines 10,000
Table 1: Overall statistics of the data volumes from different medical knowledge resources.
We integrate multiple types of retrievers to optimize our retrieval performance, incluing BGE-Large-EN-v1.5222https://huggingface.co/BAAI/bge-large-en-v1.5 for dense retrieval and SPLADE-v3333https://huggingface.co/naver/splade-v3 for sparse retrieval. We then consolidate the documents retrieved through both methods into a unified collection that contains $n$ documents, $\mathcal{D}=\left\{d_{i}\right\}_{i=1}^{n}=\mathcal{D}^{D}\cup\mathcal{D}^{S}$, where $\mathcal{D}^{D}$ comprises documents obtained by dense retrieval, and $\mathcal{D}^{S}$ represents those acquired via sparse retrieval.
@@ 3.2.2 Evidence Reranker
For the document collections retrieved in Section 3.2.1, we employ a coarse-to-fine strategy for reranking. we initially utilize the BGE-Reranker-v2-M3444https://huggingface.co/BAAI/bge-reranker-v2-m3 to rerank $\mathcal{D}=\left\{d_{i}\right\}_{i=1}^{n}$ based on their semantic relevance to the current query at a coarse granularity, returning the top $k$ documents, denoted as $\mathcal{S}^{c}=\left\{d_{i}\right\}_{i=1}^{k}$. At this stage, the total length of these $k$ documents significantly exceeds the context limit of the model. Then we conduct a fine-grained reranking that integrates three distinct criteria for the current $k$ documents to get $\mathcal{S}^{f}=\left\{d_{i}\right\}_{i=1}^{k}$, enabling the model to provide answers based on the most effective retrieved documents within the limited length of the context window. The fine-grained reranking score can be formulated as:
@@
where $f_{h}\left(x\right)$, $f_{u}\left(x\right)$ and $f_{g}\left(x\right)$ refer to scores of each refined evaluation criteria, while $\alpha$ is the non-negative hyper-parameter for weight controlling.
Hierarchy of Evidence  This criterion serves dual purposes: scoring and conflict filtering. The recalled documents that have undergone coarse-grained reranking are first categorized according to their evidence levels. Subsequently, contents of these documents are analyzed for conflicting facts, where the documents that contain conflicting facts and have lower evidence ratings are filtered out. Formally, each retrieved document $d_{i}\in\mathcal{S}^{c}$ is associated with an integer evidence level $e$, where $e\in\left\{x\in\mathbb{Z}\mid 1\leq x\leq 9\right\}$, with 1 indicating the highest level of credibility:
@@
General Document Category  It corresponds to the mapping of general natural language question types discussed in Section 3.1. Different categories of questions desire distinct answer structures. For instance, questions regarding procedural steps are ideally answered by documents that describe processes rather than those that define concepts. To address this, we categorize the retrieved documents into 16 document types, denoted as $C$, as outlined in Table 5, and score them based on their alignment with the response type preferred by the original query. The scoring function is defined as follows:
@@ 3.3 Evidence Applying
CoT Generator  This module constructs a chain-of-thought reasoning process based on the original medical query and the retrieved evidence documents. It serves dual functions: (1) Component of Query Reformulation: It contributes to the subsequent query reformulation process, facilitating the retrieval of evidence documents relevant to the question. (2) Few-Shot Learning Instance: It provides few-shot examples for the LLM physician (target model) intended for downstream task evaluating, demonstrating how to analyze the retrieved evidence and address medical queries effectively.
@@ 3.4 Effect Assessment
Our assessment of evidence encompasses two key aspects: the accuracy of the CoT generator in responding to queries and the stability of evidence document retrieval across different stages, described in Algorithm 2. This dual focus ensures that when evaluating the LLM physician (target model) on medical tasks, the retrieved evidence documents provided are both effective and robust.
@@ 4 Experiments and Results
4.1 Experimental Setup
Model Details. We employ the latest open-source LLMs from LLaMA Touvron et al. (2023); Dubey et al. (2024) and Qwen Yang et al. (2024) series as our target language models for evaluation. We assessed models including Qwen2.5-7B, LLaMA3.1-8B, LLaMA2-13B, Qwen2.5-14B, Qwen2.5-32B and LLaMA3.1-70B, scaling from 7B to 70B. The default setting of context window for our main experiments is 4K, with an in-depth scaling analysis presented in Section 5.
Datasets. We have selected five medical datasets including PubMedQA Jin et al. (2019), MedQA-USMLE, MedQA-MCMLE Jin et al. (2020), MedMCQA Pal et al. (2022), and MMLU-Med Hendrycks et al. (2021), and use accuracy as evaluation metrics. More details about the datasets can be found in Section E.1.
@@ 4.1 Experimental Setup
Med-R2 85.52 85.01 74.48 78.01 84.65 81.53
Table 2: Comparison of Med-R2 with baselines. The best and second best results are in bold and underlined.
Implementation. We constructed the medical knowledge corpus by establishing the FAISS vector library Johnson et al. (2019). Experiments related to model training were conducted based on full-parameter fine-tuning, during which we utilized a learning rate scheduler featuring linear warm-up and cosine decay, peaking at a learning rate of 2e-5, alongside a warmup ratio of 0.03, a weight decay of 0.0 and a batch size of 128 for 3 epochs. We conducted all training and evaluation experiments on NVIDIA RTX H800 with 80G memory.

### v4

# Med-R2: Crafting Trustworthy LLM Physicians via Retrieval and Reasoning of Evidence-Based Medicine

@@ Abstract
Large Language Models (LLMs) have exhibited remarkable capabilities in clinical scenarios. Despite their potential, existing works face challenges when applying LLMs to medical settings. Strategies relying on training with medical datasets are highly cost-intensive and may suffer from outdated training data. Leveraging external knowledge bases is a suitable alternative, yet it faces obstacles such as limited retrieval precision and poor effectiveness in answer extraction. These issues collectively prevent LLMs from demonstrating the expected level of proficiency in mastering medical expertise. To address these challenges, we introduce Med-R2, a novel LLM physician framework that adheres to the Evidence-Based Medicine (EBM) process, efficiently integrating retrieval mechanisms as well as the selection and reasoning processes of evidence, thereby enhancing the problem-solving capabilities of LLMs in healthcare scenarios and fostering a trustworthy LLM physician. Our comprehensive experiments indicate that Med-R2 achieves an improvement of 14.74% over vanilla RAG methods and even a 3.32% enhancement compared to fine-tuning strategies, without incurring additional training costs. Furthermore, we find that our LLaMA3.1-70B + Med-R2 surpasses frontier models, including GPT-4o, Claude3.5-Sonnet and DeepSeek-V3 by 1.22%, 5.33% and 2.80%. Code and datasets supporting this study are included in our supplementary materials.
@@ 1 Introduction
Large Language Models (LLMs) have emerged as pivotal tools in the medical domain, redefining the contours of healthcare practice [1, 52, 43]. Their ability to process and understand vast amounts of unstructured medical data positions them at the forefront of medical research [8, 34], clinical decision-making [13, 27], and patient care [5, 45]. However, despite their remarkable capabilities, LLMs encounter several challenges when applied specifically to healthcare settings:
C1: Inefficiency in Knowledge Acquisition. Existing approaches predominantly rely on pretraining and fine-tuning LLMs with datasets from the medical domain, the process of which is inherently compute-intensive and resource-demanding, especially as model sizes increase [42, 32]. Moreover, training with outdated datasets can result in a lack of highly specialized expertise, which can lead to suboptimal clinical recommendations or misinform healthcare professionals [15, 25].
C2: Limited Precision for Medical Retrieval. Compared to domain-specific training, Retrieval-Augmented Generation (RAG) systems provide a cost-efficient solution, leveraging the external knowledge base to enhance content generation [28]. The retrieval quality is critical in RAG systems, where inaccuracy or misinformation can heavily influence the effectiveness of LLMs’ augmentation. While efforts [46, 19, 31] have been made to improve the retrieval precision, they neglect the specific and highly professional nature of medical knowledge, where tailored retrieval enhancement for distinct medical scenarios remains insufficiently explored.
@@
C3: Low Effectiveness in Answer Extraction. Considering the constraints imposed by the models’ context window length, it is essential to critically appraise the retrieved medical evidence for its validity, impact, as well as applicability, and integrate the most pertinent ones with existing clinical expertise for problem-solving [6, 11]. Nonetheless, current studies fail to develop targeted answer extraction methods tailored for healthcare scenarios, where the nuanced evaluation of evidence hierarchies and intricate reasoning are required [40].
@@
  • Challenges in Medical Scenarios. Through conducting a quantitative analysis of strategies aimed at enhancing models’ medical capabilities, we underscore the challenges prevalent in healthcare scenarios, including high computational consumption as well as poor efficiency in knowledge retrieval and extraction (Figure 1 and Section 4).
  • LLM Physician Framework. We present Med-R2, a novel LLM physician framework that integrates the principles of evidence-based medicine (EBM) for clinical problem-solving with outstanding retrieval and reasoning, bolstering the performance of models within medical contexts (Section 3).
  • Performance and Effectiveness. Our extensive experiments indicate that Med-R2 achieves a 14.74% improvement over the vanilla RAG methods, and even a 3.32% enhancement compared to the fine-tuning strategies without additional training expenses. Moreover, LLaMA3.1-70B + Med-R2 surpasses frontier models for medical problem-solving, achieving average improvements over GPT-4o, Claude3.5-Sonnet and DeepSeek-V3 by 1.22%, 5.33% and 2.80% (Section 4 and Section 5).
2 Related Work and Discussions
Evidence-Based Medicine (EBM)  EBM refers to the application of the best available research to healthcare, which requires evidence integration with clinical expertise and patient values [40, 11, 39]. Clinical questions can be categorized into several types, including diagnosis, therapy, prognosis, etiology, prevention, cost, etc. [6]. Each category intersects with EBM principles by emphasizing the systematic collection, evaluation, and application of the best retrieved evidence to inform medical decision-making [12], detailed in Appendix A.
@@ 2 Related Work and Discussions
LLMs for Medical Domain  As the application of LLMs expands, their deployment in the medical domain has become a widely discussed topic [8, 54]. Recent studies have concentrated on the direct use of real or synthetic medical data for the pretraining or fine-tuning of LLMs [43, 41]. Prominent open-source milestones include ChatDoctor [29] which integrates real-world doctor-patient communication data for training, PMC-LLaMA [47] pretrained on 4.9 million medical literature records, and MEDITRON [7], a scaling series of medical pretrained models. However, such extensive training can be computationally intensive. In contrast, Retrieval-Augmented Generation (RAG) systems offer a more efficient alternative, achieving comparable results with reduced training costs and enhancing the model’s precision in locating and leveraging knowledge.
Retrieval-Augmented Generation (RAG)  The concept of RAG [28] was introduced as a powerful framework for integrating external knowledge into natural language generation tasks, enhancing the accuracy and relevance of generated outputs across various domains [10, 55, 18]. In the medical field, RAG has been widely used to improve LLMs’ analytical performances by utilizing external medical knowledge from sources such as medical papers, textbooks, guidelines, and entries [23, 53, 49, 50]. However, while there have been efforts dedicated to optimizing the individual components of RAG pipelines [46, 19, 4, 20], research that integrates the unique characteristics and requirements of the medical domain remains in its infancy. In this study, we incorporate the principles of Evidence-Based Medicine (EBM) into medical RAG systems to better address the special demands of healthcare.
@@ 3.1 Question Formulation
The clinical question formulation stage consists of two components: query classifier and query reformulator. The query classification encompasses two dimensions: Evidence-Based Medicine (EBM) categories and general natural language question types. Specifically, we have delineated six distinct EBM categories and twelve general question categories, as illustrated in Figure 2. The application of this classification scheme to the MedMCQA dataset yielded the categorized results presented in Figure 3. For the classification task, we have manually annotated 100 samples to fine-tune Qwen2.5-72B-Instruct [^1] https://huggingface.co/Qwen/Qwen2.5-72B-Instruct, and employed the trained model as our classifier. We perform domain-specific reformulations of the original queries based on their respective classes of the EBM categories to align with the professional context. Meanwhile, the general question categories are utilized as one of the criteria to rerank retrieved documents, thereby prioritizing those that best match the current query’s intent and document type preferences described in Figure 8. Further details regarding the method and implementation can be found in Section B.1.
@@ 3.2.1 Evidence Retriever
Knowledge Corpus  To construct a more comprehensive medical knowledge base adaptable to diverse healthcare scenarios, we have amassed a collection of medical data to build our retrieval corpus. The final medical knowledge base employed for retrieval comprises four distinct types of resources: academic papers, entries, books, and guidelines, with details of the data sources and statistics depicted in Section B.2.1 and Table 1. For resources with extensive content such as academic papers, books, and guidelines, we first perform content segmentation with a threshold set at 10,000 tokens. We prioritize dividing the content based on natural chapters. If natural chapters cannot be identified or exceed the threshold, we resort to truncation according to the predefined limit.
Table 1: Overall statistics of different medical knowledge resources.
Source Type #Volume #Tokens / Doc
Max Min Mean Medium
Academic Papers 600,000 10,643 279 3,820 3,097
Entries 470,000 6,538 56 1,387 1,962
Books 10,000 15,384 524 4,083 4,319
Guidelines 10,000 4,981 74 1,100 1,778
Total 1,090,000 15,384 56 2,748 2,219
We integrate multiple types of retrievers to optimize our retrieval performance, incluing BGE-Large-EN-v1.5222https://huggingface.co/BAAI/bge-large-en-v1.5 for dense retrieval and SPLADE-v3333https://huggingface.co/naver/splade-v3 for sparse retrieval. We further discuss the rationale for choosing hybrid dense and sparse retrieval in Section B.2.2. We then consolidate the documents retrieved through both methods into a unified collection that contains $n$ documents, $\mathcal{D}=\left\{d_{i}\right\}_{i=1}^{n}=\mathcal{D}^{D}\cup\mathcal{D}^{S}$, where $\mathcal{D}^{D}$ comprises documents obtained by dense retrieval, and $\mathcal{D}^{S}$ represents those acquired via sparse retrieval.
@@ 3.2.2 Evidence Reranker
For the document collections retrieved in Section 3.2.1, we employ a coarse-to-fine strategy for reranking. we initially utilize the BGE-Reranker-v2-M3444https://huggingface.co/BAAI/bge-reranker-v2-m3 to rerank $\mathcal{D}=\left\{d_{i}\right\}_{i=1}^{n}$ based on their semantic relevance to the current query at a coarse granularity, returning the top $k$ documents [^5] During the previous retrieval phase, documents of all retrieval types are subjected to a unified ranking, within which the top-$k$ documents are selected for further processing., denoted as $\mathcal{S}^{c}=\left\{d_{i}\right\}_{i=1}^{k}$. At this stage, the total length of these $k$ documents significantly exceeds the context limit of the model. Then we conduct a fine-grained reranking that integrates three distinct criteria for the current $k$ documents to get $\mathcal{S}^{f}=\left\{d_{i}\right\}_{i=1}^{k}$, enabling the model to provide answers based on the most effective retrieved documents within the limited length of the context window. The fine-grained reranking score can be formulated as:
@@
where $f_{h}\left(x\right)$, $f_{u}\left(x\right)$ and $f_{g}\left(x\right)$ refer to scores of each refined evaluation criteria, while $\alpha$ is the non-negative hyper-parameter for weight controlling. The detailed implementation and explanation for the derivation of Equation 1 can be found in Section B.3.
Hierarchy of Evidence  Medical knowledge encompasses facts and theories that are not always consistent, and sometimes even contradictory. This criterion serves dual purposes: scoring and conflict filtering. The recalled documents that have undergone coarse-grained reranking are first categorized according to their evidence levels. Subsequently, contents of these documents are analyzed for conflicting facts, where the documents that contain conflicting facts and have lower evidence ratings are filtered out. Formally, each retrieved document $d_{i}\in\mathcal{S}^{c}$ is associated with an integer evidence level $e$, where $e\in\left\{x\in\mathbb{Z}\mid 1\leq x\leq 9\right\}$, with 1 indicating the highest level of credibility:
@@
By applying the evidence assessment criteria, the retrieved documents are categorized into multiple evidence levels, and professional analysis and integration of evidence at different levels of authority are conducted, which prevents misjudgments caused by information clutter.
@@
General Document Category  It corresponds to the mapping of general natural language question types discussed in Section 3.1. Different categories of questions desire distinct answer structures. For instance, questions regarding procedural steps are ideally answered by documents that describe processes rather than those that define concepts. To address this, we categorize the retrieved documents into 16 document types, denoted as $C$, as outlined in Table 6, and score them based on their alignment with the response type preferred by the original query. The scoring function is defined as:
@@ 3.3 Evidence Applying
CoT Generator  This module constructs a chain-of-thought reasoning process based on the original medical query and the retrieved evidence documents [^6] When ablating the integration of CoT sequences at the onset in Section 5, the CoT sequences are curated solely from the original medical query during the initial iteration, excluding the retrieved evidence documents.. It serves dual functions: (1) Component of Query Reformulation: It contributes to the subsequent query reformulation process, facilitating the retrieval of evidence documents relevant to the question. In Section 5, we also clarify the rationale behind not incorporating models’ generated CoT sequences into the query reformulator for medical knowledge retrieval at the initial iteration. (2) Few-Shot Learning Instance: It provides few-shot examples for the LLM physician (target model) intended for downstream task evaluating, demonstrating how to analyze the retrieved evidence, identify the key information, and address medical queries effectively.
@@ 3.4 Effect Assessment
Our assessment of evidence is centered around two key aspects: the accuracy of the chain-of-thought (CoT) generator in responding to queries and the stability of evidence document retrieval throughout different stages. Algorithm 2 delineates the evidence evaluation process and the iterative retrieval loop for the query corpus within the pipeline, where we determine the number of iterations for optimization termination based on our assessment of these two factors. This dual focus ensures that when evaluating the LLM physician (target model) on medical tasks, the retrieved evidence documents provided are both effective and robust.
@@ 4 Experiments and Results
Table 2: Comparison of Med-R2 with baselines. The best and second best are in bold and underlined.
@@
Frontier Models
GPT-4o 84.95 79.28 74.60 77.20 84.44 80.09
Claude3.5-Sonnet 83.20 72.54 68.80 76.40 83.90 76.97
DeepSeek-V3 80.93 77.82 74.30 73.60 87.67 78.86
Open-Sourced Medical Models
PMC-LLaMA-7B 38.65 35.30 32.40 52.26 26.87 37.10
MEDITRON-7B 46.76 42.78 36.83 58.32 43.60 45.66
PMC-LLaMA-13B 40.38 33.07 37.55 54.63 46.85 42.50
MEDITRON-70B 55.05 52.21 53.50 72.84 71.45 61.01
Open-Sourced Base Models
@@
Med-R2 86.37 84.58 73.36 78.24 82.82 81.07
4.1 Experimental Setup
Model Details. We employ the latest open-sourced LLMs from LLaMA [44, 9] and Qwen [51] series as our target language models for evaluation. We assessed models including Qwen2.5-7B, LLaMA3.1-8B, LLaMA2-13B, Qwen2.5-14B, Qwen2.5-32B and LLaMA3.1-70B, scaling from 7B to 70B. The default setting of context window for our main experiments is 4K, with an in-depth scaling analysis presented in Section 5.
Datasets. We have selected five medical datasets including PubMedQA [22], MedQA-USMLE, MedQA-MCMLE [21], MedMCQA [35], and MMLU-Med [16], and use accuracy as evaluation metrics. More details about the datasets can be found in Section C.1.
Implementation. We constructed the medical knowledge corpus by establishing the FAISS vector library [24]. Experiments related to model training were conducted based on full-parameter fine-tuning, during which we utilized a learning rate scheduler featuring linear warm-up and cosine decay, peaking at a learning rate of 2e-5, alongside a warmup ratio of 0.03, a weight decay of 0.0 and a batch size of 128 for 3 epochs. We conducted all training and evaluation experiments on NVIDIA RTX H800 GPUs with 80G memory.